Opened 5 weeks ago

Last modified 2 days ago

#633 new flight processing

GB18/66, flight day 178/2018, Sefton Coast

Reported by: dac Owned by:
Priority: immediate Milestone:
Component: Processing: general Keywords:
Cc: Other processors:


Data location: ~arsf/arsf_data/2018/flight_data/uk/GB18_66-2018_178_Sefton_Coast

Data arrived from NERC-ARF via hard drive on 13/07/2018

Scientific objective: Mapping and monitoring Dune fields on the Sefton Coast.

PI: Paul Aplin
BAS Project Code: D066

Due to lack of LiDAR PI requested SfM processing to generate DSM.

The following issues were identified during unpacking:

  • Missing darkframes for owl capture data.


  • Fenix (requested, flown)
  • OWL (not requested, flown)
  • Phase One (requested, flown)

Change History (15)

comment:1 Changed 5 weeks ago by dac


GPS stop time for line 1 was listed as xx:xx:xx.x, replaced with 13:43:43, calculated using FPS and number of lines.

comment:2 Changed 4 weeks ago by dac

Digital Camera

Converted raw images to tiffs.

comment:3 Changed 3 weeks ago by asm

  • Summary changed from GB18/66, flight day 178/2011, Sefton Coast to GB18/66, flight day 178/2018, Sefton Coast

comment:4 Changed 3 weeks ago by asm

Navigation Processing

Started. Downoladed basestation BLAP from ​OS National GPS Network. Used navigation files downloaded by IpasTC from closest basestation. Basestation verification:
Lat: 53 46 36.94825
Lon: -3 02 05.82891
Ell. Height 66.193 m

Good results on navigation with position separation smaller than 0.2 m (good for Fenix Pixel size) and ambiguity fix 2. Files copied acros. Navigation Processing finished.

comment:5 Changed 3 weeks ago by asm

Last edited 3 weeks ago by asm (previous) (diff)

comment:6 Changed 3 weeks ago by asm


Wrote PI on 27/07/2018 to ask him if they deployed their own basestation data as they have done in the past so we can use it for processing. Further processing is for now on hold until we have a reply.

comment:7 Changed 2 weeks ago by dac

Hyperspectral Processing

Ran through with existing .sol file to check navigation sync. Doesn't look to be any problems. SCT value of 2.0 works for most lines apart from 3 where 1.0 looks good. Match well with OS vector data.

Holding off further processing until we hear back about GPS basestation.

comment:8 Changed 9 days ago by asm

Hyperspectral Processing

Found SCT values:

Flightline FENIX
1 2
2 2
3 1
4 2
5 2
6 2
7 2
8 2
9 2
10 2
11 1
12 1
13 2

Navigation was going backwards for line 11, using -force flag in aplnav solved the issue. This is likely to be related to the persistent problem we are having with the fps of the instrument this season. The lines overlap OK with very small offsets so can't correct the fps manually as it is within acceptance values. Will create a delivery and add a note on readme.

comment:9 Changed 4 days ago by dac

Navigation processing

Heard back from PI on 06/08/2018. GPS basestation wasn't set up for flight, continuing processing using OS RINEX data.

comment:10 Changed 3 days ago by asm

Hyperspectral Processing

Delivery and readme created. Hyperspectral procesing finished, ready for DC.

comment:11 Changed 3 days ago by dac

Hyperspectral Delivery Check

Starting delivery check.

comment:12 Changed 3 days ago by dac

Hyperspectral Delivery Check

  • There was 'logsheet.pdf' in the root directory as well as the correctly named one in logsheet. I checked the one in the directory logsheet which was newer. logsheet.pdf needs to be removed (unless this is the correct one)
  • Logsheet is missing PI
  • In readme, quality and issues with data section suggest following changes to text:
    • During this season, the framerate of the instrument has been not responding accordingly and could not be set correctly. This led to a small "stretching" of the mapped data and inaccurate geocorrection of the Fenix data in other flights > During this season the framerate of the instrument has sometimes being recording incorrectly, this has led to a small stretching of the mapped data and inaccurage coecorrection for these flights.
    • via QGIS (delete).
  • removed fodis directory
  • Registration accuracy looks good, it could be improved with a better DEM (which is what we normally recommend).
  • Colours in the screenshots look a bit off, and there are lots of black patches in the middle around the dunes, data look OK though so should be fine.
  • Spectral look good when compared to Py6S simulations

After minor updates have been made to logsheet and readme ready to deliver.

comment:13 Changed 3 days ago by asm

Hyperspectral Delivery

Made changes suggested to the logsheet and README files. Will zip mapped files.

comment:14 Changed 3 days ago by asm

Hyperspectral Delivery

Mapped files were zipped correctly. PI selected hard drive as method of delivery so started copying files across.

comment:15 Changed 2 days ago by asm

Hyperspectral Delivery

Delivered via FTP slot 5 and hard drive sent to PI. Notification sent on 17/08/2018. Finished

Note: See TracTickets for help on using tickets.